# FDF: Flexible Decoupled Framework for Time Series Forecasting with Conditional Denoising and Polynomial Modeling

### v1

@@ C.3. Ablation experiments
  • FDF1: Removes the conditional information.
  • FDF2: Replaces the PTM with a simple linear layer.
  • FDF3: Removes instance normalization during both the training and inference phases.
  • FDF4: Removes the conditional information and replaces PTM with a linear layer.
  • FDF5: Removes instance normalization alignment.
The results show that each component in FDF impacts overall performance to varying degrees, particularly in modeling complex temporal patterns. Removing the condition information in FDF1 leads to a slight increase in MSE on ETTh2 (from 0.3393 to 0.3404), indicating that while this module helps integrate external information, the model remains resilient without it, though with reduced flexibility. More significant degradation occurs when replacing the PTM with a simple linear layer (FDF2), especially on ETTm2, where MSE increases from 0.2303 to 0.2450, emphasizing the importance of the non-linear trend modeling provided by the PTM. The removal of instance normalization (FDF3) results in the sharpest performance drop, particularly on ETTm2, where MSE rises to 0.2459. This underscores the critical role of instance normalization in stabilizing training and preventing overfitting on datasets with significant temporal variations. Removing both the condition information and the PTM in FDF4 leads to further performance degradation, showing that these components play complementary roles in improving model performance. Finally, eliminating instance normalization alignment during training (FDF5) yields mixed results. On ETTh2, a slight improvement suggests that aligning normalization parameters may hinder the model’s ability to adapt to complex dynamics, while on ETTm2, where temporal dependencies dominate, the effect is minimal. Overall, the ablation study validates FDF’s modular design in improving adaptability and accuracy across diverse time series forecasting tasks.

### v2

@@ C.3. Ablation experiments
  • FDFw/o ${}_{\text{w/o CR}}$: Removes the conditional information.
[…]
  • FDFw/o ${}_{\text{w/o INR}}$: Removes instance normalization during both the training and inference phases.
  • FDFw/o ${}_{\text{w/o CR+PLR}}$: Removes the conditional information and replaces PTM with a linear layer.
  • FDFw/o ${}_{\text{w/o INA}}$: Removes instance normalization alignment.
The results show that each component in FDF impacts overall performance to varying degrees, particularly in modeling complex temporal patterns. Removing the condition information in FDFCR leads to a slight increase in MSE on ETTh2 (from 0.3393 to 0.3404), indicating that while this module helps integrate external information, the model remains resilient without it, though with reduced flexibility. More significant degradation occurs when replacing the PTM with a simple linear layer (FDFPLR), especially on ETTm2, where MSE increases from 0.2303 to 0.2450, emphasizing the importance of the non-linear trend modeling provided by the PTM. The removal of instance normalization (FDF3) results in the sharpest performance drop, particularly on ETTm2, where MSE rises to 0.2459. This underscores the critical role of instance normalization in stabilizing training and preventing overfitting on datasets with significant temporal variations. Removing both the condition information and the PTM in FDF4 leads to further performance degradation, showing that these components play complementary roles in improving model performance. Finally, eliminating instance normalization alignment during training (FDF5) yields mixed results. On ETTh2, a slight improvement suggests that aligning normalization parameters may hinder the model’s ability to adapt to complex dynamics, while on ETTm2, where temporal dependencies dominate, the effect is minimal. Overall, the ablation study validates FDF’s modular design in improving adaptability and accuracy across diverse time series forecasting tasks.